# Collectively enhanced ground-state cooling in subwavelength atomic arrays

### v2

@@ I Introduction
Previous studies have addressed this problem leading to different and often contradicting results. For large one-dimensional atomic arrays Palmer and Beige [29] demonstrated that dipole interaction can lead to a $30\%$ reduction of the final motional temperature. The scope of their work is however limited to a minimum interatomic separation, as their description breaks down in the presence of perfectly dark states. Smaller separations were considered by Shahmoon et al [30, 31] and Robicheaux and co-workers [32, 33], demonstrating that collective dipole interactions generally lead to excess motional heating in 2D arrays when the time scale of the motion is respectively much longer and much shorter than the scattering dynamics. To further complicate the picture, Wang et al [34] demonstrated that a single target emitter in free space can be cooled to a slightly lower temperature if it is dipole coupled to a small surrounding ring of emitters, albeit only for specific configurations that reduce coherent-dipole exchange between atoms. However, this strategy cannot be directly generalized to enhance the cooling of multiple emitters simultaneously.
@@
The remainder of the article is structured as follows. In Sec. II, we introduce the model and derive the effective master equation for the atomic motion. After reviewing the cooling of independent (i. e., non-interacting) atoms in Sec. III, we present, in Sec. IV, the first main results of our work: the collective cooling of two dipole-interacting atoms. This simple example illustrates the main conceptual contributions of this study. We then generalize these results to arrays of $N$ atoms in Sec. V. Our conclusions are summarized in Sec. VI. The article is complemented by several appendices; in particular, Appendix B details the derivation of the effective master equation for atomic motion, which incorporates collective effects.
@@ II System and Model
We consider an ensemble of $N$ two-level atoms with mass $M$ and resonant dipole transition (|$\ket{g}\leftrightarrow\ket{e}$) at a frequency $\omega_{0}=ck_{0}=2\pi c/\lambda_{0}$ and linewidth $\gamma_{0}$. The atoms are harmonically trapped at positions $\mathbf{r}_{0j}$ and are driven by an external monochromatic laser at frequency $\omega_{L}=\omega_{0}+\Delta$ [Fig. 1(a)]. Their internal degrees of freedom are described by the lowering operators $\hat{\sigma}_{j}\equiv\ket{g_{j}}\bra{e_{j}}$ |, while their motional degrees of freedom are described by the center-of-mass momentum $\hat{\mathbf{p}}_{j}$ and position $\hat{\mathbf{r}}_{j}=\mathbf{r}_{0j}+\delta\hat{\mathbf{r}}_{j}$ operators, where $\delta\hat{\mathbf{r}}_{j}$ denotes the displacement from the trap center. Following this description, we will refer to internal excitation of the atoms as spin excitations. Throughout this work, we focus on ordered atomic arrays with subwavelength interatomic spacing $d<\lambda_{0}$. Nevertheless the results of this section apply more broadly to ensembles of trapped atoms with arbitrary geometries and interatomic separations.
@@
$\begin{split}\hat{H}_{\text{A}}\equiv&\sum_{j}\left[{\frac{\hat{\mathbf{p}}_{j}}{2M}+\frac{M\nu_{j}^{2}}{2}(\hat{\mathbf{r}}_{j}-\mathbf{r}_{0j})^{2}}\right]-\hbar\Delta\sum_{j}\hat{\sigma}^{\dagger}_{j}\hat{\sigma}_{j}\\ &+\hbar\sum_{i,j}G(\hat{\mathbf{r}}_{i}-\hat{\mathbf{r}}_{j})\hat{\sigma}^{\dagger}_{i}\hat{\sigma}_{j},\end{split}$ (2)
@@
where $J(\mathbf{r})$ and $\gamma(\mathbf{r})$ represent the coherent and dissipative contributions, respectively. At the origin, we define $G(\mathbf{0})=-i\gamma_{0}/2$. The explicit form of the dipole-coupling rate Eq. (3) depends on the electromagnetic Green’s tensor that describes how photon propagates and mediate interactions between the atoms. In the following we consider atoms in free space for which the explicit form of Eq. (3) is given in Appendix A.
@@ II.1 Lamb-Dicke approximation
where we defined the bosonic annihilation (creation) operator $\hat{b}_{j}$ ($\hat{b}^{\dagger}_{j}$) which creates (destroys) a motional quantum (phonon) at site $j$. The first term is the mechanical energy of the atoms in the trap, while the second term describes the carrier excitation of the laser, |$\ket{g_{j}}\leftrightarrow\ket{e_{j}}$, which does not change the phonon number. The last term reads
@@
Here we defined the non-hermitian Hamiltonian
@@
where $\Omega_{j}^{\prime\prime}\equiv k_{0}^{-2}\partial^{2}\Omega(\mathbf{r})/\partial^{2}r_{\alpha}|_{\mathbf{r}_{0j}}$, and $G^{\prime\prime}_{ij}$ is defined above. The population recycling terms, respectively the second and last term in Eq. (13), capture the diffusion of the center-of-mass motion induced by photon scattering, now occurring in a correlated fashion due to dipole-dipole interactions. Equations (13) and (14) are obtained from the second order Lamb-Dicke term of Eq. (1), except for the term that, as a result of normal ordering of the phonon operators, is independent of the motional degrees of freedom. We have subtracted this contribution from (13) and (14) and included it in Eq. (8). It describes effects of the finite spatial width of the ground-state wavefunctions in the harmonic traps, that effectively smears out the dipole-dipole interactions over the extent of the atomic motion. This modifies the linear optical response of the system as compared to the case of non-moving atoms, as we shall now illustrate.
The system’s linear optical response is obtained from the single-excitation spectrum of Eq. (10). We define the eigenstates |$\ket{\phi_{\lambda}}=\sum_{j}\phi_{\lambda,j}\hat{\sigma}_{j}^{\dagger}\ket{g}$ such that $H_{S}\ket{\phi_{\lambda}}=\epsilon_{\lambda}\ket{\phi_{\lambda}}$. The complex eigenvalues $\epsilon_{\lambda}=-\Delta+J_{\lambda}-i\gamma_{\lambda}/2$ describe the collective energy shifts $J_{\lambda}$ and collective linewidth $\gamma_{\lambda}$ of the transition |$\ket{g}^{\otimes N}\leftrightarrow\ket{\phi_{\lambda}}$. As a consequence of interference between the field emitted by nearby atoms in the array, the linewedith $\gamma_{\lambda}$ is strongly modified leading to bright ($\gamma_{\lambda}\gg\gamma_{0}$) and dark ($\gamma_{\lambda}\ll\gamma_{0}$) collective resonances. While for an array of pinned atoms, collective dark states have a vanishing linewidth in the limit of $N\rightarrow\infty$, this is not the case for Eq. (10). The finite extent of the atomic ground state motional wavefunction broaden the collective linewdith such that [35, 36]
@@
This key feature prevents the emergence of perfectly dark states. We remark that the broadening in Eq. (15) is solely due to vacuum fluctuations. Additional corrections arise from finite temperature contribution and from dipole-induced forces. These terms, respectively included in Eq. (11) and Eq. (13), are of the same order in $\eta^{2}$ and contribute to the broadening of the collective linewidth (see Appendix B).
@@ II.2 Effective master equation for the motional dynamics
$\begin{split}\frac{d}{dt}\hat{\rho}&=-i\Big{[}\sum_{i,j=1}^{N}V_{ij}e^{\mathrm{i}(\nu_{i}-\nu_{j})t}\hat{b}^{\dagger}_{i}\hat{b}_{j},\hat{\rho}\Big{]}\\ &+\sum_{i,j=1}^{N}\!\!\mathcal{R}_{ij}^{-}e^{\mathrm{i}(\nu_{i}-\nu_{j})t}\left(\hat{b}_{j}\hat{\rho}\hat{b}_{i}^{\dagger}-\frac{1}{2}\{\hat{b}_{i}^{\dagger}\hat{b}_{j},\hat{\rho}\}\right)\\ &+\sum_{i,j=1}^{N}\!\!\mathcal{R}_{ij}^{+}e^{-\mathrm{i}(\nu_{i}-\nu_{j})t}\left(\hat{b}_{j}^{\dagger}\hat{\rho}\hat{b}_{i}-\frac{1}{2}\{\hat{b}_{i}\hat{b}_{j}^{\dagger},\hat{\rho}\}\right).\end{split}$ (16)
@@ III Sideband cooling of independent atoms
$\bar{n}_{\mathrm{ind}}=\langle\hat{b}^{\dagger}\hat{b}\rangle=\frac{\mathcal{R}^{+}}{\mathcal{R}^{-}-\mathcal{R}^{+}},$ (18)
@@
±$\mathcal{R}^{\pm}=\frac{\eta^{2}|\Omega^{\prime}|^{2}\gamma_{0}}{(-\Delta\pm\nu)^{2}+(\gamma_{0}/2)^{2}}+\frac{\eta^{2}|\gamma_{0}^{\prime\prime}||\Omega|^{2}}{\Delta^{2}+(\gamma_{0}/2)^{2}}.$ (19)
@@ IV Collective sideband cooling of two atoms
Figure 3: (a) Two harmonically trapped atoms separated by $d<\lambda_{0}$ are coupled by coherent (dissipative) $J_{12}$ ($\gamma_{12}$) dipole interactions and are driven by an external laser along the direction of motion ($z$). (b) Motional sidebands: a cooling and a heating sideband is associated to the symmetric (superradiant) and to the antisymmetric (subradiant) transition. The dashed grey lines indicate the sidebands of non-interacting atoms. (c) For $\nu_{1}\approx\nu_{2}\approx\nu$, the center-of-mass (relative) motion of both atoms is independently cooled via the symmetric (antisymmetric) cooling sideband to phonon numbers higher (lower) than that of non-interacting atoms. (d) For larger trap frequency difference, both atoms are cooled to lower temperatures via the cooling antisymmetric sideband. (e) Normalized average steady state phonon number, $\bar{n}/\bar{n}_{\textit{ind}}$, as a function of frequency difference $\delta\nu$ and drive strength $\Omega$ for $d=0.2\lambda_{0}$, $\nu=20\gamma_{0}$ and $\Delta\approx-\nu-J_{12}$, obtained by solving the Lamb-Dicke master equation Eq. (7). The dashed black lines correspond to the two conditions in Eq. (32), which separate the three cooling regimes. The dash-dotted white line corresponds parameters used to compute the critical cooling rate in Fig. 4(b) for $d=0.2\lambda_{0}$.
The linear response of two dipole-coupled atoms is characterized by two distinctive resonances associated with the excitation of the ground state |$\ket{gg}$ to either the symmetric state |$\ket{S}=(\ket{ge}+\ket{eg})/\sqrt{2}$ or the antisymmetric state |$\ket{A}=(\ket{ge}-\ket{eg})/\sqrt{2}$ [41, 42]. The transition to the symmetric (antisymmetric) states are characterized by a resonance frequency $J_{S}\equiv\omega_{0}+J_{12}+\eta^{2}J^{\prime\prime}_{12}$ ($J_{A}\equiv\omega_{0}-J_{12}-\eta^{2}J^{\prime\prime}_{12}$) and by a collective linewidth $\gamma_{S}=\gamma_{0}+\gamma_{12}+\eta^{2}\gamma^{\prime\prime}_{12}$ ($\gamma_{A}=\gamma_{0}-\gamma_{12}-\eta^{2}\gamma^{\prime\prime}_{12}$). At short distances ($d<\lambda_{0}/2$), the symmetric state becomes superradiant with an enhanced decay rate $\gamma_{S}>\gamma_{0}$, while the antisymmetric state becomes subradiant with a narrowed linewidth $\gamma_{A}<\gamma_{0}$. In the limit $d\ll\lambda_{0}$, the dipole-dipole interaction can be expanded in powers of the small parameter $k_{0}d$. The coherent coupling scales as $J_{12}\propto(k_{0}d)^{-3}$, while the dissipative terms $\gamma_{12}$ and $\gamma_{12}^{\prime\prime}$ remain finite. In this limit, the symmetric state decays at $\gamma_{S}=2\gamma_{0}+\mathcal{O}(k_{0}^{2}d^{2})$, whereas the antisymmetric state becomes strongly subradiant, with decay rate
@@
[…]
@@
Figure 4: Sideband cooling of two interacting atoms in the array cooling regime for $\eta=0.02$. (a) Average phonon number $\bar{n}$ as a function of average trap frequency $\bar{\nu}$ for various atomic separations $d$, with $\delta\nu=10^{-3}\gamma_{0}$. Solid line are obtained from Eq. (16) while diamond markers are obtained by simulating Eq. (7). (i-iii) Scaling of $\bar{n}$ with $d$ for various $\bar{\nu}$. The black dashed lines correspond to the analytical estimates given by Eqs. (34)-(37). (b-c) Critical cooling rate $\Gamma_{c}$ as a function of $\delta\nu$ for various $d$ and for (b) $\bar{\nu}=20\gamma_{0}$ and (c) $\bar{\nu}=0.25\gamma_{0}$. For the case $d=0.2\lambda_{0}$ shown in panel (b), the driving strength corresponding to the critical cooling rate is indicated by the white dash-dotted line in Fig. 3(e). In all panels, the dashed grey lines correspond to the results for non-interacting atoms, while the solid lines and diamond markers are respectively obtained by solving the effective master equation (16) and the full model including the spin dynamics. For the numerical simulation in panel (a) we set $\Omega/\gamma_{0}=10^{-3}$.
@@ IV.1 Single-mode cooling
At the optimal detuning, the dissipative motional coupling associated to cooling, $\mathcal{R}_{12}^{-}$, is larger than that associated to heating, $\mathcal{R}_{12}^{+}$, and over the coherent motional coupling $V_{12}$  [^1] For example, this can be easily in the regime $\bar{\nu}\gg\gamma_{0}$, where $R_{12}^{+}\propto\eta^{2}\Omega^{2}\gamma_{0}/\bar{\nu}^{2}$ and $V_{12}\propto\eta^{2}\Omega^{2}/\bar{\nu}$.. Thus, Eq. (27) can be interpreted as the condition under which the motional degrees of freedom of both atoms are resonantly coupled. Then, the effective master equation for the atomic motion becomes diagonal when expressed in terms of the center-of-mass and relative motional modes, defined by the operators $\hat{b}_{S}=(\hat{b}_{1}+\hat{b}_{2})/\sqrt{2}$ and $\hat{b}_{A}=(\hat{b}_{1}-\hat{b}_{2})/\sqrt{2}$, respectively. The problem thus reduces to two independent instances of sideband cooling, where the center-of-mass and relative motion are respectively cooled by the symmetric and antisymmetric spin states at rates $\mathcal{R}_{S}^{\pm}$ and $\mathcal{R}_{A}^{\pm}$, as illustrated in Fig. 3(c).
@@
Compared to the steady-state phonon number for non-interacting atoms in the resolved sideband regime, given by Eq. (20), the occupations $\bar{n}_{S}$ and $\bar{n}_{A}$ are modified by the decay properties of the collective spin states. At short distances $d<\lambda_{0}/2$, $\gamma_{A}<\gamma_{0}$ and hence $\bar{n}_{A}<\bar{n}_{\text{ind}}^{\text{(sc)}}$. In contrast, the symmetric state is superradiant, leading to $\bar{n}_{S}>\bar{n}_{\text{ind}}^{(\text{sc})}$.
While collective interactions enable enhanced cooling of one collective motional mode (in this case, the relative motion), the average phonon number of the system is necessarily larger than that of two non-interacting atoms,

### v3

@@ I Introduction
Previous studies have addressed this problem leading to different and often contradicting results. The case of two ions was studied by Vogt et al. [29] unveiling regimes of enhanced cooling due to both subradiance and interference between different scattering processes. It is however unclear whether this mechanism generalizes to many atoms. For large one-dimensional atomic arrays, Palmer and Beige [30] demonstrated that dipole interaction can lead to a $30\%$ reduction of the final motional temperature. The scope of their work is however limited to a minimum interatomic separation because their description breaks down in the presence of perfectly dark states. Smaller separations were considered by Shahmoon et al. [31, 32] and Robicheaux and co-workers [33, 34], demonstrating that collective dipole interactions generally lead to excess motional heating in two-dimensional (2D) arrays when the timescale of the motion is respectively much longer and much shorter than the scattering dynamics. To further complicate the picture, Wang et al [35] demonstrated that a single target emitter in free space can be cooled to a slightly lower temperature if it is dipole coupled to a small surrounding ring of emitters, albeit only for specific configurations that reduce coherent-dipole exchange between atoms. However, this strategy cannot be directly generalized to enhance the cooling of multiple emitters simultaneously.
@@
The remainder of the article is structured as follows. In Sec. II, we introduce the model and derive the effective master equation for the atomic motion. After reviewing the cooling of independent (i. e., non-interacting) atoms in Sec. III, we present, in Sec. IV, the first main results of our work: the collective cooling of two dipole-interacting atoms. This simple example illustrates the main conceptual contributions of this study. We then generalize these results to arrays of $N$ atoms in Sec. V. Our conclusions are summarized in Sec. VI. The article is complemented by several appendixes; in particular, Appendix B details the derivation of the effective master equation for atomic motion, which incorporates collective effects.
@@ II System and Model
[…]
@@
$\begin{split}\hat{H}_{\text{A}}\equiv&\sum_{j}\left[{\frac{\hat{\mathbf{p}}^{2}_{j}}{2M}+\frac{M\nu_{j}^{2}}{2}(\hat{\mathbf{r}}_{j}-\mathbf{r}_{0j})^{2}}\right]-\hbar\Delta\sum_{j}\hat{\sigma}^{\dagger}_{j}\hat{\sigma}_{j}\\ &+\hbar\sum_{i,j}G(\hat{\mathbf{r}}_{i}-\hat{\mathbf{r}}_{j})\hat{\sigma}^{\dagger}_{i}\hat{\sigma}_{j},\end{split}$ (2)
@@
where $J(\mathbf{r})$ and $\gamma(\mathbf{r})$ represent the coherent and dissipative contributions, respectively. At the origin, we define $G(\mathbf{0})=-i\gamma_{0}/2$. The explicit form of the dipole-coupling rate (3) depends on the electromagnetic Green’s tensor that describes how the photon propagates and mediate interactions between the atoms. In the following, we consider atoms in free space for which the explicit form of Eq. (3) is given in Appendix A.
@@ II.1 Lamb-Dicke approximation
[…]
@@
Here we defined the nonhermitian Hamiltonian
@@
where $\Omega_{j}^{\prime\prime}\equiv k_{0}^{-2}\partial^{2}\Omega(\mathbf{r})/\partial^{2}r_{\alpha}|_{\mathbf{r}_{0j}}$, and $G^{\prime\prime}_{ij}$ is defined above. The population recycling terms, respectively the second and last term in Eq. (13), capture the diffusion of the center-of-mass motion induced by photon scattering, now occurring in a correlated fashion due to dipole-dipole interactions. Equations (13) and (14) are obtained from the second order Lamb-Dicke term of Eq. (1), except for the term that, as a result of normal ordering of the phonon operators, is independent of the motional degrees of freedom. We have subtracted this contribution from (13) and (14) and included it in Eq. (8). It describes effects of the finite spatial width of the ground-state wavefunctions in the harmonic traps, that effectively smears out the dipole-dipole interactions over the extent of the atomic motion. This modifies the linear optical response of the system as compared with the case of nonmoving atoms, as we shall now illustrate.
[…]
@@
This key feature prevents the emergence of perfectly dark states. We remark that the broadening in Eq. (15) is solely due to vacuum fluctuations. Additional corrections arise from finite temperature contribution and from dipole-induced forces. These terms, respectively included in Eqs. (11) and (13), are of the same order in $\eta^{2}$ and contribute to the broadening of the collective linewidth (see Appendix B).
@@ II.2 Effective master equation for the motional dynamics
[…]
@@ III Sideband cooling of independent atoms
$\bar{n}_{\mathrm{ind}}=\langle\hat{b}^{\dagger}\hat{b}\rangle$$\frac{\mathcal{R}^{+}}{\mathcal{R}^{-}-\mathcal{R}^{+}}$, (18)
@@
$\mathcal{R}^{\pm}=\frac{\eta^{2}|\Omega^{\prime}|^{2}\gamma_{0}}{(-\Delta\pm\nu)^{2}+(\gamma_{0}/2)^{2}}+\frac{\eta^{2}|\gamma_{0}^{\prime\prime}||\Omega|^{2}}{\Delta^{2}+(\gamma_{0}/2)^{2}}.$ (19)
@@ IV Collective sideband cooling of two atoms
Figure 3: (a) Two harmonically trapped atoms separated by $d<\lambda_{0}$ are coupled by coherent (dissipative) $J_{12}$ ($\gamma_{12}$) dipole interactions and are driven by an external laser along the direction of motion ($z$). (b) Motional sidebands: a cooling and a heating sideband is associated to the symmetric (superradiant) and to the antisymmetric (subradiant) transition. The dashed gray lines indicate the sidebands of noninteracting atoms. (c) For $\nu_{1}\approx\nu_{2}\approx\nu$, the center-of-mass (relative) motion of both atoms is independently cooled via the symmetric (antisymmetric) cooling sideband to phonon numbers higher (lower) than that of noninteracting atoms. (d) For larger trap frequency difference, both atoms are cooled to lower temperatures via the cooling antisymmetric sideband. (e) Normalized average steady state phonon number, $\bar{n}/\bar{n}_{\textit{ind}}$, as a function of frequency difference $\delta\nu$ and drive strength $\Omega$ for $d=0.2\lambda_{0}$, $\nu=20\gamma_{0}$ and $\Delta\approx-\nu-J_{12}$, obtained by solving the Lamb-Dicke master equation (7). The dashed black lines correspond to the two conditions in Eq. (32), which separate the three cooling regimes. The dash-dotted white line corresponds parameters used to compute the critical cooling rate in Fig. 4(b) for $d=0.2\lambda_{0}$.
[…]
@@
$\hat{H}_{L}^{(1)}\!=\!\frac{i\hbar\Omega}{\sqrt{2}}\!\left[|S\rangle\langle gg|\left(\!\hat{R}_{1}+\hat{R}_{2}\!\right)+|A\rangle\langle gg|\left(\!\hat{R}_{1}-\hat{R}_{2}\!\right)\right]\!+\text{h.c.}$ (23)
@@
Figure 4: Sideband cooling of two interacting atoms in the array cooling regime for $\eta=0.02$. (a) Average phonon number $\bar{n}$ as a function of average trap frequency $\bar{\nu}$ for various atomic separations $d$, with $\delta\nu=10^{-3}\gamma_{0}$. Solid line are obtained from Eq. (16) while diamond markers are obtained by simulating Eq. (7). (i-iii) Scaling of $\bar{n}$ with $d$ for various $\bar{\nu}$. The black dashed lines correspond to the analytical estimates given by Eqs. (34)-(37). (b-c) Critical cooling rate $\Gamma_{c}$ as a function of $\delta\nu$ for various $d$ and for (b) $\bar{\nu}=20\gamma_{0}$ and (c) $\bar{\nu}=0.25\gamma_{0}$. For the case $d=0.2\lambda_{0}$ shown in panel (b), the driving strength corresponding to the critical cooling rate is indicated by the white dash-dotted line in Fig. 3(e). In all panels, the dashed gray lines correspond to the results for noninteracting atoms, while the solid lines and diamond markers are respectively obtained by solving the effective master equation (16) and the full model including the spin dynamics. For the numerical simulation in panel (a) we set $\Omega/\gamma_{0}=10^{-3}$.
@@ IV.1 Single-mode cooling
At the optimal detuning, the dissipative motional coupling associated to cooling, $\mathcal{R}_{12}^{-}$, is larger than that associated to heating, $\mathcal{R}_{12}^{+}$, and over the coherent motional coupling $V_{12}$  [^1] For example, this can be easily in the regime $\bar{\nu}\gg\gamma_{0}$, where $R_{12}^{+}\propto\eta^{2}\Omega^{2}\gamma_{0}/\bar{\nu}^{2}$ and $V_{12}\propto\eta^{2}\Omega^{2}/\bar{\nu}$.. Thus, Eq. (27) can be interpreted as the condition under which the motional degrees of freedom of both atoms are resonantly coupled. Then, the effective master equation for the atomic motion becomes diagonal when expressed in terms of the center of mass and relative motional modes, defined by the operators $\hat{b}_{S}=(\hat{b}_{1}+\hat{b}_{2})/\sqrt{2}$ and $\hat{b}_{A}=(\hat{b}_{1}-\hat{b}_{2})/\sqrt{2}$, respectively. The problem thus reduces to two independent instances of sideband cooling, where the center-of-mass and relative motion are respectively cooled by the symmetric and antisymmetric spin states at rates $\mathcal{R}_{S}^{\pm}$ and $\mathcal{R}_{A}^{\pm}$, as illustrated in Fig. 3(c).
@@
Compared with the steady-state phonon number for noninteracting atoms in the resolved sideband regime, given by Eq. (20), the occupations $\bar{n}_{S}$ and $\bar{n}_{A}$ are modified by the decay properties of the collective spin states. At short distances $d<\lambda_{0}/2$, $\gamma_{A}<\gamma_{0}$ and hence $\bar{n}_{A}<\bar{n}_{\text{ind}}^{\text{(sc)}}$. In contrast, the symmetric state is superradiant, leading to $\bar{n}_{S}>\bar{n}_{\text{ind}}^{(\text{sc})}$.
While collective interactions enable enhanced cooling of one collective motional mode (in this case, the relative motion), the average phonon number of the system is necessarily larger than that of two noninteracting atoms,